OpenStudy (anonymous):

(-1)/((2)sqrt(21-x))=-1/5 SOLVE for x

OpenStudy (anonymous):

this problem is similar to your last problem. Square on both sides. 1/(4(21-x)) = 1/25 4(21-x)=25 21-x=25/4 x=21-25/4 do the rest so that you will learn.